0.02x+0.0175(15,000-x)=277.5
Answer by tutorcecilia(2152)

0.02x+0.0175(15,000-x)=277.5 [Simplify - distribute]
0.02x + 262.5 - 0.0175x = 277.5 [Combine like-terms]
0.0025x + 262.5 = 277.5 [Solve for x]
0.0025x + 262.5 - 262.5 = 277.5 - 262.5
0.0025x = 15
0.0025x/0.0025=15/0.0025
x = 6000
.
check by plugging (x=6000) back into the original equation